4 times an integer is 9 more than 3 times the next consecutive integer. What are the two integers?
Answer by jim_thompson5910(35256)


4 times an integer is 9 more than 3 times the next consecutive integer

translates to

4x = 3(x+1) + 9

solve for x...

4x = 3(x+1) + 9

4x = 3x+3 + 9

4x = 3x+12

4x - 3x = 12

x = 12

x+1 = 13

The two integers are 12 and 13.